arricake:

A swimming pool has a length of 30 feet, a width of 15 feet, and a depth of 4 feet. How much water can the swimming pool hold?

carmelle:

To find the volume, you have to multiply the length, width, and depth. \[30~\times 15~\times~4~=~?\]

Krissy525:

It would be L*W*D L=30 W=15 D=4 30*15*4=?

achaf0644:

L=length=30ft w=width=15ft d=depth=4 so LxWxD=A(answer) meaning 30x15x4=A time to solve: 30x10=300+30x5=150+300=450 400x4=1200+50x4=200+1200=1400ft so 30x15x4=1400 feet Hope this helps! Have a great day!

hiiiii2221:

30 * 15 * 4 = 1800 ft³.

hiiiii2221:

@achaf0644 wrote:
L=length=30ft w=width=15ft d=depth=4 so LxWxD=A(answer) meaning 30x15x4=A time to solve: 30x10=300+30x5=150+300=450 400x4=1200+50x4=200+1200=1400ft so 30x15x4=1400 feet Hope this helps! Have a great day!
you are incorrect, you calculated it wrong

achaf0644:

*L=length=30ft w=width=15ft d=depth=4 so LxWxD=A(answer) meaning 30x15x4=A time to solve: 30x10=300+30x5=150+300=450 400x4=1600+50x4=200+1200=1800ft so 30x15x4=1800 feet Hope this helps! Have a great day! sorry i was wrong the first time. i hope this redo makes up for it. thank you @hiiiii2221 for noticing that. :)

hiiiii2221:

@achaf0644 wrote:
*L=length=30ft w=width=15ft d=depth=4 so LxWxD=A(answer) meaning 30x15x4=A time to solve: 30x10=300+30x5=150+300=450 400x4=1600+50x4=200+1200=1800ft so 30x15x4=1800 feet Hope this helps! Have a great day! sorry i was wrong the first time. i hope this redo makes up for it. thank you @hiiiii2221 for noticing that. :)
No problem